Write the simplest polynomial function with the given zeros (-2, 2, 4)
Found 2 solutions by solver91311, Edwin McCravy:
Answer by solver91311(24713)



If a number is a zero of a polynomial, then the binomial is a factor of the polynomial. The simplest polynomial function with the given zeros is the polynomial function with the three factors that correspond to the three given zeros. Since there are three zeros, the polynomial function must be of the third degree.

John


Answer by Edwin McCravy(20063)

Write the simplest polynomial function with the given zeros (-2, 2, 4)

The simplest polynomial function with real zeros 

z1, z2, z3, ... zn 

is found by starting with 

P(x) = (x-z1)(x-z2)...(x-zn)

and multiplying it out.

So the simplest polynomial function with xeros -2, 2, and 4 is

P(x) = [x-(-2)](x-2)(x-4)

P(x) = [x+2](x-2)(x-4)

P(x) = (x+2)(x-2)(x-4)

P(x) = (x2-2x+2x-4)(x-4)

P(x) = (x2-4)(x-4)

P(x) = x3-4x2-4x+16

Edwin